OpenStudy (anonymous):

Help please ! Ill give medals for the person who can help me.. Find the area of an equiangular triangle circumscribed in a circle with a radius of 203

OpenStudy (anonymous):

@n648c788

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@Lyrae

OpenStudy (anonymous):

Look it up on google. Thats what i usually do .

OpenStudy (anonymous):

wut..

OpenStudy (anonymous):

@n648c788

OpenStudy (anonymous):

Sorry I am still new to this site, so I did not see your message 1.So Formula r=2*h/3, radius=203 units 203 units=2*h/3 so Height= 304.5 units 2. Then for base or Side Formula, Base or side=2(h)/sqrt3 2(304.5 units)/sqrt3 Base= 351.61units 3.Finally A=1/2bh A=1/2(351.61 units* 304.5 units) A=53532.62 units2

OpenStudy (anonymous):

Message me for any questions! I will get back to you this time